OpenStudy (angiedoesschool2275):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price.What was the price of the flowers in September?

Vocaloid (vocaloid):

October price = (1+0.2)*(September price) 1.50 = (1.2)*(September price) solve for September price

OpenStudy (angiedoesschool2275):

What are the steps to answering the problem? don't understand your message Vocaloid...

Vocaloid (vocaloid):

"The October price was a 20% increase of the September price." means that October price = (1.2)*(September price) October price = 1.50, therefore 1.50 = (1.2)*(September price) so, what is the September price?

Vocaloid (vocaloid):

hint: divide both sides of the equation by 1.2

OpenStudy (angiedoesschool2275):

So basically it's just 1.50 x 1.2?

Vocaloid (vocaloid):

1.50 = (1.2)*(September price) 1.50/1.2 = September price

OpenStudy (angiedoesschool2275):

So the September price is 1.8?

Vocaloid (vocaloid):

please look carefully 1.50/1.2 = September price = ?

Vocaloid (vocaloid):

/ means divide

OpenStudy (angiedoesschool2275):

1.25?

Vocaloid (vocaloid):

yes

OpenStudy (angiedoesschool2275):

So the total answer is 1.25? no more steps?

Vocaloid (vocaloid):

yes